# Geometry Aware Operator Transformer As An Efficient And Accurate Neural Surrogate For PDEs On Arbitrary Domains

### v3

@@ 1 Introduction
To answer this question, we have performed extensive ablation studies in SM Sec.F to observe that i) the MAGNO encoder/decoder is clearly superior to message-passing based encoders/decoders, ii) choosing a regular equispaced latent point cloud performs significantly better than either downsampling on the original point cloud or using a projected low-dimensional regular grid, iii) GAOT is highly robust to the size of its latent grid, iv) a time-derivative marching strategy, i.e, setting $\gamma=1,\delta=\tau$ in (7) is superior to other choices of $\gamma,\delta$, v) using a statistical geometric embedding performs significantly better than either not using additional geometric information or using a pointnet to process geometric information vi) incorporating multiscale features in the MAGNO encoder/decoder provides a significant gain in accuracy when compared to using just a single scale GNO encoder/decoder as in GINO and vii) The power of GAOT does not just stem from its VIT processor, but also from its MAGNO encoder/decoder (SM E.5), acting in tandem. These results justify the choices that we have made in designing GAOT and selecting the relevant model components, while also revealing how these innovative features underpin GAOT’s accuracy.
However, as argued before, this accuracy might come at the price of computational inefficiency. But, as we have demonstrated above, GAOT is also the most efficient model and does not have to pay the accuracy-efficiency trade-off. The reasons behind this boil down to the tricks used in efficiently implementing GAOT, which are discussed at length in SM E.2 and E.1.
@@ 4 Discussion
Summary. We present GAOT, a new neural operator for learning PDE solutions on arbitrary domains. It is based on a novel multiscale GNO encoder/decoder, combined with geometric embeddings that convey statistical information about the local domain geometry, and a (vision) transformer based processor. The model is designed to handle any point cloud input and provide the output at any query point in the underlying domain. Several innovative strategies have been used to make the implementation of GAOT computationally efficient and scalable. We test GAOT on a large number of challenging datasets for a variety of time-dependent and time-independent PDEs over diverse two-dimensional domain geometries to find that GAOT is significantly more accurate, robust and computationally efficient in terms of training throughput and inference latency, over a large set of baselines. We further demonstrate the potential of GAOT by presenting its SOTA performance on three large-scale three-dimensional datasets of industrial simulations in the automobile and aerospace sectors. These results demonstrate that GAOT can be a powerful and scalable neural operator with wide-spread applications. They also showcase the main advantage with an efficient and accurate neural operator such as GAOT, i.e, its inference time is many orders of magnitude faster than the runtime of a classical numerical PDE solver. We quantify this speedup in SM E.11 to find that GAOT can be anywhere between 4 to 9 orders of magnitude faster to run than classical PDE solvers, while retaining accuracy.
Related Work. As discussed before, there are 3 broad classes of models for learning PDEs on arbitrary domains namely i) end-to-end message-passing based frameworks exemplified here with RIGNO, which significantly improves upon models such as (multiscale) MeshGraphNets MGN ; MSMGN ii) Transformer based frameworks such as Transolver wu2024transolver , GNOT hao2023gnot and UPT UPT and iii) frameworks, based on GNO encoders/decoders and FNO processors as in GINO li2023geometryinformed . GAOT differs from all these approaches by a) not using graph-based message passing, b) only employing transformers in the processor c) using a transformer, rather than FNO as a processor and significantly augmenting the GNO encoder/decoder by multiscale features, attention based-quadrature and geometry embeddings. It is precisely these choices, along with a highly efficient implementation, that allows GAOT to significantly surpass GINO, RIGNO, and Transolver in both accuracy and efficiency.
Limitations and Extensions. GAOT’s excellent scalability and ability to generalize very well in a transfer learning setting (Fig. 3(c)) showcase its potential to serve as the backbone of foundation models for PDEs, extending models such as Poseidon herde2024poseidon and DPOT hao2024dpot to arbitrary domains. Physics-informed loss functions can be added to GAOT to enable it to act as a physics-informed neural operator as in li2024physics . We also plan to apply GAOT to downstream tasks such as UQ LMR1 , inverse problems, NIO and PDE constrained optimization ISMO to further test its abilities. Finally, theoretical results for GAOT, such as universal approximation, will be considered in future work.
@@ Acknowledgements
The contribution of Siddhartha Mishra to this work was supported in part by the DOE SEA-CROGS project (DE-SC-0023191).
@@ D.1 Poisson-C-Sines
This dataset contains solutions to the two-dimensional Poisson equation with a circular domain. The Poisson equation is a fundamental linear elliptic partial differential equation (PDE) given by Eq. D.2. The dataset represents the mapping from the source term $f$ to the solution $u$ using the solution operator $\mathcal{G}^{\dagger}:f\mapsto u$. The source term is defined as:
@@
where r=0.5r=-0.5 and $K=16$. The coefficients $a_{ij}$ are sampled i.i.d. uniformly from [1,1][-1,1] to generate the dataset. The solution $u$ is computed on a circular domain with zero Dirichlet boundary conditions. The dataset is generated using a finite element method (FEM) on a triangular mesh in a circular domain. The mesh is generated using the Delaunay algorithm with 16431 points and 32441 elements.
@@ E.8 Comparison of Neural Field with UPT
E.9 Transfer Learning
@@ E.10 Training Randomness
E.11 Runtime Comparison between GAOT and Classical Solvers

### v4

@@ 1 Introduction
To answer this question, we have performed extensive ablation studies in SM Sec.F to observe that i) the MAGNO encoder/decoder is clearly superior to message-passing based encoders/decoders, ii) choosing a regular equispaced latent point cloud performs significantly better than either downsampling on the original point cloud or using a projected low-dimensional regular grid, iii) GAOT is highly robust to the size of its latent grid, iv) a time-derivative marching strategy, i.e, setting $\gamma=1,\delta=\tau$ in (7) is superior to other choices of $\gamma,\delta$, v) using a statistical geometric embedding performs significantly better than either not using additional geometric information or using a pointnet to process geometric information vi) incorporating multiscale features in the MAGNO encoder/decoder provides a significant gain in accuracy when compared to using just a single scale GNO encoder/decoder as in GINO and vii) The power of GAOT does not just stem from its VIT processor, but also from its MAGNO encoder/decoder (SM E.5), acting in tandem. These results justify the choices that we have made in designing GAOT and selecting the relevant model components, while also revealing how these innovative features underpin GAOT’s accuracy. However, as argued before, this accuracy might come at the price of computational inefficiency. But, as we have demonstrated above, GAOT is also the most efficient model and does not have to pay the accuracy-efficiency trade-off. The reasons behind this boil down to the tricks used in efficiently implementing GAOT, which are discussed at length in SM E.2 and E.1.
@@ 4 Discussion
Summary. We present GAOT, a new neural operator for learning PDE solutions on arbitrary domains. It is based on a novel multiscale GNO encoder/decoder, combined with geometric embeddings that convey statistical information about the local domain geometry, and a (vision) transformer based processor. The model is designed to handle any point cloud input and provide the output at any query point in the underlying domain. Several innovative strategies have been used to make the implementation of GAOT computationally efficient and scalable. We test GAOT on a large number of challenging datasets for a variety of time-dependent and time-independent PDEs over diverse two-dimensional domain geometries to find that GAOT is significantly more accurate, robust and computationally efficient in terms of training throughput and inference latency, over a large set of baselines. We further demonstrate the potential of GAOT by presenting its SOTA performance on three large-scale three-dimensional datasets of industrial simulations in the automobile and aerospace sectors. These results demonstrate that GAOT can be a powerful and scalable neural operator with wide-spread applications. They also showcase the main advantage with an efficient and accurate neural operator such as GAOT, i.e, its inference time is many orders of magnitude faster than the runtime of a classical numerical PDE solver. We quantify this speedup in SM E.12 to find that GAOT can be anywhere between 4 to 9 orders of magnitude faster to run than classical PDE solvers, while retaining accuracy.
Related Work. As discussed before, there are 3 broad classes of models for learning PDEs on arbitrary domains namely i) end-to-end message-passing based frameworks exemplified here with RIGNO, which significantly improves upon models such as (multiscale) MeshGraphNets Pfaff et al. (2021); Fortunato et al. (2022) ii) Transformer based frameworks such as Transolver Wu et al. (2024), GNOT Hao et al. (2023) and UPT Alkin et al. (2024) and iii) frameworks, based on GNO encoders/decoders and FNO processors as in GINO Li et al. (2023b). GAOT differs from all these approaches by a) not using graph-based message passing, b) only employing transformers in the processor c) using a transformer, rather than FNO as a processor and significantly augmenting the GNO encoder/decoder by multiscale features, attention based-quadrature and geometry embeddings. It is precisely these choices, along with a highly efficient implementation, that allows GAOT to significantly surpass GINO, RIGNO, and Transolver in both accuracy and efficiency.
@@ Acknowledgements
The contribution of Siddhartha Mishra to this work was supported in part by the DOE SEA-CROGS project (DE-SC-0023191). The authors thank Mohamed Elrefaie and Prof. Faez Ahmed (MIT) for their inputs regarding the Carbench datasets and results.
@@ D.1 Poisson-C-Sines
This dataset contains solutions to the two-dimensional Poisson equation with a circular domain. The Poisson equation is a fundamental linear elliptic partial differential equation (PDE) given by Eq. D.2. Unlike standard benchmarks that map point-wise evaluations of the source term to the solution, this dataset defines the operator learning task within the context of the finite element discretization. Specifically, we aim to learn the mapping from the projection coefficients of the source term $f$ onto the finite element basis functions to the nodal solution values $u$.
The continuous source term is constructed via a spectral expansion:
@@
where r=0.5r=-0.5 and $K=16$. The coefficients $a_{ij}$ are sampled i.i.d. uniformly from [1,1][-1,1]. To generate the input data, this source term is projected onto the finite element test space spanned by the nodal basis functions $\{\phi_{k}\}$. The input to the model is the resulting load vector $\mathbf{b}$, where each entry corresponds to the integral $b_{k}=\int_{\Omega}f(x,y)\phi_{k}(x,y)d\Omega$. Consequently, the operator learning task is defined as $\mathcal{G}^{\dagger}:\mathbf{b}\mapsto\mathbf{u}$, representing the inverse of the discretized stiffness operator. The mesh is generated using the Delaunay algorithm with 16,431 points and 32,441 elements, and zero Dirichlet boundary conditions are enforced during the assembly.
@@ E.8 Comparison of Neural Field with UPT
Table E.8: Quantitative comparison on the DrivAerNet++ dataset. The performance on both subsampled point clouds (10k points) and full-resolution meshes (487k points) is taken from the CarBench paper [15]. While baseline models struggle with the high-resolution input, GAOT is trained directly on the full-resolution mesh, enabling it to capture fine-grained geometric details. We report results for a representative unseen test sample (E_S_WW_WM_648) and the dataset-wide average across all 1154 test samples.
Model MAE ($\mathbf{m}^{2}/\mathbf{s}^{2}$) RMSE ($\mathbf{m}^{2}/\mathbf{s}^{2}$) Rel L2 $\mathbf{R}^{2}_{\textbf{test}}$
10k Full 10k Full 10k Full 10k Full
PointNet 32.2 32.5 62.4 63.3 0.395 0.403 0.757 0.747
NeuralOperator 25.9 26.3 51.2 54.4 0.324 0.347 0.837 0.813
PointMAE 23.0 24.7 46.2 52.0 0.292 0.331 0.867 0.829
PointNetLarge 22.0 25.1 42.1 55.3 0.266 0.352 0.890 0.807
RegDGCNN 16.5 19.9 32.4 46.1 0.205 0.293 0.935 0.866
PointTransformer 16.2 20.6 29.5 47.0 0.187 0.299 0.946 0.861
TransolverLarge 12.2 18.5 25.5 47.2 0.161 0.300 0.959 0.860
TripNet 13.6 19.8 25.1 44.0 0.158 0.297 0.961 0.863
Transolver++ 12.3 18.3 24.6 46.0 0.156 0.293 0.962 0.867
Transolver 11.7 17.7 23.7 46.1 0.150 0.294 0.965 0.866
AB-UPT 10.2 17.2 21.5 45.7 0.136 0.291 0.971 0.869
GAOT (Sample 648) 11.4 22.8 0.145 0.967
GAOT (Avg. 1154 samples) 12.9 23.9 0.157 0.958
E.9 Training at full vs. low Resolution for the DrivAernet++ Dataset.
Given these results, we turn our attention to the recent investigation of the question of whether training at full resolution confers any benefits over training at low resolution and simply inferring at full resolution. This question was recently explored in the context of the industrial scale DrivAernet++ dataset within the Carbench framework in [15]. In their study, the authors of [15] trained various AI surrogates at a subsampled resolution of $10K$ points from the full surface point cloud of $487$ K points. Then, they tested the models on an unseen sample at the low-resolution of $10K$ points as well as the full resolution of $487K$ points. To ensure uniformity, they interpolated the results from testing at low-resolution to the full resolution. We reproduce their results verbatim in Table E.8 where all the rows except the last two are identical to Table 3 of [15]. As seen from the Table, the performance of models trained at low-resolution degrades severely when interpolated to the full resolution. In fact, all models suffer from this issue and models such as Transolver and AB-UPT [2] which perform very well at low-resolution suffer from so much degradation in accuracy that their performance is at par with the models such as RegDGCNN, which were also inaccurate at low-resolution.
On the other hand, we trained GAOT on the full resolution and present the results of the single test sample as well as the average for all test samples in the last two rows of Table E.8 to observe that GOAT, trained at full resolution is twice more accurate than all the other models. Moreover, the performance of GAOT trained on full resolution is on par with the models that been trained and even tested on the low-resolution. Thus, this experiment further indicates that training at full resolution, if possible, is the best strategy for yielding accurate neural surrogates.
Finally, we would like to point out that models such as AB-UPT [2] also possess the neural field property and can be tested at full resolution, when trained at low resolution, using the neural field property, rather than just interpolating to high resolution. However, as shown in Table 3, we recall that the neural field property does not necessarily translate to performance retention and one can expect a possible degradation of accuracy, even with this paradigm, while training at low resolutions. Nevertheless, this issue needs to investigated further.
E.10 Transfer Learning
@@ E.11 Training Randomness
E.12 Runtime Comparison between GAOT and Classical Solvers